# EDiT: A Local-SGD-Based Efficient Distributed Training Method for Large Language Models

### v1

@@ Abstract
Distributed training methods are crucial for large language models (LLMs). However, existing distributed training methods often suffer from communication bottlenecks, stragglers, and limited elasticity. Local SGD methods have been proposed to address these issues, but their effectiveness remains limited to small-scale training due to additional memory overhead and lack of concerns on efficiency and stability. To tackle these issues, we propose EDiT, an innovative Efficient Distributed Training method that combines a tailored Local SGD approach with model sharding techniques to enhance large-scale training efficiency. EDiT performs layer-wise parameter synchronization during forward pass, reducing communication and memory overhead and enabling the overlap of computation and communication. Besides, EDiT employs a pseudo gradient penalty strategy to suppress loss spikes, which ensures training stability and improve performance. Additionally, we introduce A-EDiT, a fully asynchronous variant of EDiT that accommodates heterogeneous clusters. Building on EDiT/A-EDiT, we conduct a series of experiments to validate large-scale asynchronous training for LLMs, accompanied by comprehensive analyses. Experimental results demonstrate the superior performance of EDiT/A-EDiT, establishing them as robust solutions for distributed LLM training in diverse computational ecosystems.
@@ 1 Introduction
Distributed training stands as the cornerstone for deep neural networks (Dean et al., 2012). With the explosive growth of model scale and data volume (Touvron et al., 2023; Bai et al., 2023), techniques such as Zero Redundancy Optimizer (ZeRO) (Rajbhandari et al., 2020) and 3D Parallelism (Narayanan et al., 2021) are precipitated to address the challenges of GPU memory overflow. These approaches rely on synchronous paradigm, which introduces significant communication overhead during the training process (Douillard et al., 2023). Besides, the synchronous paradigm also introduces the straggler problem, where faster workers are idle waiting for the slower ones to catch up. This issue is particularly prevalent in large/heterogeneous clusters (Lian et al., 2018). Lastly, in resource-constrained clusters, there is a compelling need for elastic training (Li et al., 2023). However, synchronous training paradigms struggle in elastic settings, where dynamic scaling of resources disrupts optimal hyperparameters and requires slow redistribution of model parameters.
These challenges have spurred significant research into distributed optimization methods. A typical method is Local Stochastic Gradient Descent (a.k.a Local SGD or Local-Update SGD) (Zhang et al., 2016), where each worker independently executes multiple local optimization steps in parallel before averaging model parameters across all workers. Subsequent studies have improved upon this foundational paradigm to improve the performance (Lin et al., 2019; Wang et al., 2019; Douillard et al., 2023). However, existing Local SGD methods are not easily applicable to the training of large language models (LLMs). These methods do not handle model sharding well, preventing their application to models larger than billions of parameters. Moreover, previous applications of Local SGD have focused on small-scale, highly curated datasets (Zhang et al., 2016; Douillard et al., 2023), making their results less transferable to LLM training that relies on vast, noisy datasets where instability may be introduced during the training process. Although current Local SGD methods can improve overall training efficiency by reducing communication frequency and diminishing the impact of random stragglers, they still struggle with the presence of consistent stragglers within heterogeneous devices (Liu et al., 2024). Additionally, because most existing Local SGD methods synchronize all parameters globally after some local optimizer updates, extra communication overhead are introduced (Sun et al., 2023). Lastly, current Local SGD methods predominantly employ a uniform averaging strategy to synchronize the parameters, failing to fully capitalize on the inherent differences in training progress across diverse workers (Douillard et al., 2023).
@@
To address these challenges, we propose a novel Efficient Distributed Training (EDiT) method for large language models. As illustrated in Figure 1, EDiT employs a hierarchical distribution strategy on a two-dimensional device mesh, where all workers are data parallel. Model parameters are fully sharded along the model shard dimension and synchronized along the model sync dimension. With the efficient communication links within the model shard groups and the low-frequency periodic synchronization strategy within the model sync groups, the impact of communication overhead and random stragglers is effectively alleviated. When synchronizing parameters, EDiT operates layer by layer during the forward pass and makes use of a prefetch strategy to overlap computation and communication, thereby reducing the additional communication and GPU memory overhead introduced by parameter synchronization. Additionally, EDiT employs a novel pseudo-gradient penalty method to address the instability caused by workers progressing differently due to the diverse large-scale corpus, and it leverages these differences among workers. Furthermore, we propose an asynchronous variant of the EDiT method named A-EDiT to deal with the consistent stragglers in heterogeneous clusters. We conducted a comprehensive evaluation of our proposed methods on LLM tasks, demonstrating its effectiveness compared to state-of-the-art methods.  [^1] The code will be released after the paper is accepted.
@@ 3.1 Overview
    During the forward pass of the $l$-th module, if the current updated step requires model synchronization, i.e., $(t*\tau+p)>t_{warm}$ and $p==0$ where $t$ is the outer step, $p$ is the inner step, $t_{warm}$ is the number of warmup steps and $\tau$ is the synchronization interval, parameters are synced in model sync groups, as outlined in lines 7 to 9 in Algorithm 1. In practice, the communication overhead is minimal due to the large synchronization interval ($\tau\gg 1$) and sharded parameters. Herein a novel pseudo gradient penalty strategy is introduced to enhance training stability that will be detailed. After that, each worker gathers the full module parameters through its model shard group for forward computations and promptly frees excess parameters to conserve memory.
@@
Different from other Local SGD methodologies that synchronize parameters after optimizer update, EDiT performs layer-wise parameter synchronization during forward pass. In practice, we normally employ a prefetch strategy that aggregates parameters for the upcoming module concurrently with ongoing computations, with which communications within model sync groups can be effectively overlapped with forward computations. In this way, EDiT further diminishes the additional communication overhead introduced by parameter synchronization.
@@ 3.2 Pseudo Gradient Penalty
In EDiT, the extra parameters and outer momentum on each worker are sharded in correspondence with the sharded parameters. Compared to previous methods that maintain full extra parameters and outer momentum on each worker, EDiT reduces the additional memory usage. Additionally, based on layer-wise synchronization and prefetch strategy, EDiT can further offload the extra parameters and outer momentum to the CPU and only transfer the corresponding layer’s data to GPU as needed, thereby minimizing memory overhead. Since the data for each layer is relatively small, the GPU-CPU data transfer can be effectively overlapped with GPU computations and GPU-GPU communications, ensuring fast parameter synchronization.
@@ A.3.1 Convergence and Generalization
In the main text, we present the performance of EDiT and other Local SGD methods in training the Llama 1B model in Figure 4 and Table 1. To demonstrate that EDiT performs consistently well across different model scales, we additionally trained Llama 350M, 3B, and 7B models using EDiT on the FineWeb-Edu dataset, each with a total of 420B tokens. The corresponding training loss, validation PPL, and evaluation results are shown in Figure 8 and Table 5. It can be observed that EDiT is robust across various model scales. Besides, to our knowledge, this is the first time to train a 7B model on a large-scale dataset with a Local SGD-related method. Although CO2 (Sun et al., 2023) also claimed that they trained a 7B model, they only used about 50B tokens and provided only the final validation PPL results. The models trained with EDiT will be open-sourced upon the acceptance of this paper.
@@ A.3.3 Scalability
Here we provide the detailed training loss curves of the Baseline and EDiT methods under different numbers of workers and distinct learning rates, which correspond to the Figure 6 in the main text.

### v2

@@ Abstract
Distributed training methods are crucial for large language models (LLMs). However, existing distributed training methods often suffer from communication bottlenecks, stragglers, and limited elasticity, particularly in heterogeneous or large-scale environments. Local SGD methods have been proposed to address these issues, but their effectiveness remains limited to small-scale training due to additional memory overhead and lack of concerns on efficiency and stability. To tackle these issues, we propose EDiT, an innovative Efficient Distributed Training method that combines a tailored Local SGD approach with model sharding techniques to enhance large-scale training efficiency. EDiT performs layer-wise parameter synchronization during forward pass, reducing communication and memory overhead and enabling overlap. Besides, EDiT employs a pseudo gradient penalty strategy to suppress loss spikes, which ensures training stability and improves performance. Additionally, we introduce A-EDiT, a fully asynchronous variant of EDiT that accommodates heterogeneous clusters. Building on EDiT/A-EDiT, we conduct a series of experiments to validate large-scale asynchronous training for LLMs, accompanied by comprehensive analyses. Experimental results demonstrate the superior performance of EDiT/A-EDiT, establishing them as robust solutions for distributed LLM training in diverse computational ecosystems. The code is available at Atorch codebase  [^1] https://github.com/intelligent-machine-learning/atorch/tree/main/atorch/local_sgd.
@@ 1 Introduction
With the explosive growth of model scale and data volume (Touvron et al., 2023; Bai et al., 2023), distributed methods (Rajbhandari et al., 2020; Narayanan et al., 2021; Dean et al., 2012) become increasingly critical for training deep neural networks. These approaches rely on synchronous paradigm, which introduces significant communication overhead during the training process (Douillard et al., 2023). Besides, the synchronous paradigm also introduces the straggler problem, where faster workers are idle waiting for the slower ones to catch up. This issue is particularly prevalent in large/heterogeneous clusters (Lian et al., 2018). Lastly, in resource-constrained clusters, there is a compelling need for elastic training (Li et al., 2023). However, synchronous training paradigms struggle in elastic settings, where dynamic scaling of resources disrupts optimal hyperparameters.
These challenges have spurred significant research into distributed optimization methods. A typical method is Local Stochastic Gradient Descent (a.k.a Local SGD or Local-Update SGD) (Zhang et al., 2016), where each worker independently executes multiple local optimization steps in parallel before averaging model parameters across all workers. Subsequent studies have improved upon this foundational paradigm to improve the performance (Lin et al., 2019; Wang et al., 2019; Douillard et al., 2023). However, existing Local SGD methods are not easily applicable to the training of large language models (LLMs). These methods do not handle model sharding well, preventing their application to models larger than billions of parameters. Moreover, they have focused on small-scale, highly curated datasets (Zhang et al., 2016; Douillard et al., 2023), making their results less transferable to LLM training that relies on vast, noisy datasets where instability may be introduced during training. Besides, although current Local SGD methods diminish the impact of random stragglers, they still struggle with the presence of consistent stragglers within heterogeneous devices (Liu et al., 2024). Additionally, in most existing Local SGD methods, parameter synchronization operations will introduce non-overlapped communication overhead (Sun et al., 2023). Lastly, current Local SGD methods predominantly employ a uniform averaging strategy to synchronize the parameters, failing to fully capitalize on the inherent differences in training progress across diverse workers (Douillard et al., 2023).
@@
To address these challenges, we propose a novel Efficient Distributed Training (EDiT) method for large language models. As illustrated in Figure 1, EDiT employs a hierarchical distribution strategy on a two-dimensional device mesh, where all workers are data parallel. Model parameters are fully sharded along the model shard dimension and synchronized along the model sync dimension. With the efficient communication links within the model shard groups and the low-frequency periodic synchronization strategy within the model sync groups, the impact of communication overhead and random stragglers is effectively alleviated. When synchronizing parameters, EDiT operates layer by layer during the forward pass and makes use of a prefetch strategy to overlap computation and communication, thereby reducing the additional communication and GPU memory overhead introduced by parameter synchronization. Additionally, EDiT employs a novel pseudo-gradient penalty method, which addresses the instability problem caused by diverse large-scale corpus and leverages the differences among workers to improve performance. Furthermore, we propose an asynchronous variant of the EDiT method named A-EDiT to deal with the consistent stragglers in heterogeneous clusters. We conducted a comprehensive evaluation of our proposed methods on LLM tasks, demonstrating its effectiveness compared to state-of-the-art methods.
@@ 3.1 Overview
    During the forward pass of the $l$-th module, if the current updated step requires model synchronization, i.e., $(t*\tau+p)>t_{warm}$ and $p==0$ where $t$ is the outer step, $p$ is the inner step, $t_{warm}$ is the number of warmup steps and $\tau$ is the synchronization interval, parameters are synced in model sync groups, as outlined in lines 7 to 9 in Algorithm 1. In practice, the communication overhead is minimal due to the large synchronization interval ($\tau\gg 1$) and sharded parameters. Herein a novel pseudo gradient penalty strategy is introduced to enhance training stability that will be detailed in Section 3.2. After that, each worker gathers the full module parameters through its model shard group for forward computations and promptly frees excess parameters to conserve memory.
@@
Different from other Local SGD methodologies that synchronize parameters before next step, EDiT performs layer-wise parameter synchronization during forward pass. In practice, we normally employ a prefetch strategy that aggregates parameters for the upcoming module concurrently with ongoing computations, with which communications within model sync groups can be effectively overlapped with forward computations. In this way, EDiT further diminishes the additional communication overhead introduced by parameter synchronization.
@@ 3.2 Pseudo Gradient Penalty
In EDiT, the extra parameters and outer momentum on each worker are sharded in correspondence with the sharded parameters. Compared to previous methods that maintain full extra parameters and outer momentum on each worker, EDiT reduces the additional memory usage. Additionally, based on layer-wise synchronization and prefetch strategy, EDiT can further offload the extra parameters and outer momentum to the CPU and only transfer the corresponding layer’s data to GPU as needed, thereby further minimizing memory overhead. Since the data for each layer is relatively small, the GPU-CPU data transfer can be effectively overlapped with GPU computations and GPU-GPU communications, ensuring fast parameter synchronization.
@@ A.3.1 Convergence and Generalization
In the main text, we present the performance of EDiT and other Local SGD methods in training the Llama 1B model in Figure 4 and Table 1. To demonstrate that EDiT performs consistently well across different model scales, we additionally trained Llama 350M, 3B, and 7B models using EDiT on the FineWeb-Edu dataset, each with a total of 420B tokens. The corresponding training loss, validation PPL, and evaluation results are shown in Figure 8 and Table 5. It can be observed that EDiT is robust across various model scales. Besides, to our knowledge, this is the first time to train a 7B model on a large-scale dataset with a Local SGD-related method. Although CO2 (Sun et al., 2023) also claimed that they trained a 7B model, they only used about 50B tokens and provided only the final validation PPL results.
@@ A.3.3 Scalability
Here we provide the detailed training loss curves of the Baseline and EDiT methods under different numbers of workers and distinct learning rates in Figure 10, which correspond to the Figure 6 in the main text.
@@ A.4 Proof of Theorem 1
$=$ $\mathbb{E}[\mathcal{L}(\bm{\psi}_{t,p})-\mathcal{L}(\bm{\psi}_{t,p+1})]+\nu\eta_{t,p}L\beta_{t}\sqrt{n}G_{\infty}(1-\nu\beta_{t})\sum_{k=0}^{p-1}\eta_{t,k}\mathbb{E}[\|\bm{g}_{t,k}\|]+\frac{L\nu^{2}\eta_{t,p}^{2}\beta_{t}^{2}{}nG_{\infty}^{2}}{2}$
$\leq$ $\mathbb{E}[\mathcal{L}(\bm{\psi}_{t,p})-\mathcal{L}(\bm{\psi}_{t,p+1})]+\nu\eta_{t,0}^{2}{}L\beta_{t}{}nG_{\infty}^{2}\tau+\frac{L\nu^{2}\eta_{t,p}^{2}\beta_{t}^{2}{}nG_{\infty}^{2}}{2}.$
Telescoping Equation 16 for $p=0$ to $\tau-1$ and $t=0$ to $T-1$, we have
$\sum_{t=0}^{T-1}\sum_{p=0}^{\tau-1}\nu\eta_{t,p}\beta_{t}\mathbb{E}[\|\nabla\mathcal{L}(\bm{\theta}_{t,p})\|^{2}]$ (17)
$\leq$ $\mathbb{E}[\mathcal{L}(\bm{\psi}_{0,0})-\mathcal{L}(\bm{\psi}_{T-1,\tau})]+\nu{}L{}nG_{\infty}^{2}\tau^{2}\sum_{t=0}^{T-1}\beta_{t}\eta_{t,0}^{2}+\frac{L\nu^{2}{}nG_{\infty}^{2}}{2}\sum_{t=0}^{T-1}\sum_{p=0}^{\tau-1}\beta_{t}^{2}\eta_{t,p}^{2}$
$\leq$ $\mathcal{L}(\bm{\theta}_{0,0})+\nu{}L{}nG_{\infty}^{2}\tau^{2}\sum_{t=0}^{T-1}\beta_{t}\eta_{t,0}^{2}+\frac{L\nu^{2}{}nG_{\infty}^{2}}{2}\sum_{t=0}^{T-1}\sum_{p=0}^{\tau-1}\beta_{t}^{2}\eta_{t,p}^{2}.$
Since from Equation 4, we have $1\leq\beta_{t}\leq\frac{\phi}{\epsilon}$. Combining with Equation 17, we can get
$\sum_{t=0}^{T-1}\sum_{p=0}^{\tau-1}\eta_{t,p}\mathbb{E}[\|\nabla\mathcal{L}(\bm{\theta}_{t,p})\|^{2}]\leq\frac{\mathcal{L}(\bm{\theta}_{0,0})}{\nu}+\frac{L{}nG_{\infty}^{2}\tau^{2}\phi}{\epsilon}\sum_{t=0}^{T-1}\eta_{t,0}^{2}+\frac{L\nu{}nG_{\infty}^{2}\phi^{2}}{2\epsilon^{2}}\sum_{t=0}^{T-1}\sum_{p=0}^{\tau-1}\eta_{t,p}^{2}.$ (18)
Since
$\sum_{t=0}^{T-1}\sum_{p=0}^{\tau-1}\eta_{t,p}$ $\geq\tau\sum_{t=0}^{T-1}\eta_{t,\tau-1}=\sqrt{\tau}\eta\sum_{t=0}^{T-1}\frac{1}{\sqrt{t+1}}$ (19)
@@
$=2\sqrt{\tau}\eta(\sqrt{T+1}-1)\geq 2\sqrt{\tau}\eta(\sqrt{T}-1),$
[…]
substituting Equation 19 into Equation 17, we have
$\leq\frac{1}{\sum_{t=0}^{T-1}\sum_{p=0}^{\tau-1}\eta_{t,p}}\left(\frac{\mathcal{L}(\bm{\theta}_{0,0})}{\nu}+\frac{L{}nG_{\infty}^{2}\tau^{2}\phi}{\epsilon}\sum_{t=0}^{T-1}\eta_{t,0}^{2}+\frac{L\nu{}nG_{\infty}^{2}\phi^{2}}{2\epsilon^{2}}\sum_{t=0}^{T-1}\sum_{p=0}^{\tau-1}\eta_{t,p}^{2}\right)$
This completes the proof. ∎